# A Fast Kernel-based Conditional Independence test with Application to Causal Discovery

### v1

@@ 4.1 Theoretical Insight
Under the null hypothesis $H_{0}:X\perp\!\!\!\!\perp Y\mid Z$, each block statistic $T^{(j)}_{v}$ is, by exactly the same argument as in [32], asymptotically a weighted sum of independent $\chi^{2}_{1}$ variables,
[…]
@@
Hence, our method inherits exactly the same null-law template as classical KCI: a positive, finite linear combination of $\chi^{2}_{1}$ variables. [^3] The formal requirements are the standard KCI assumptions (characteristic kernels, boundedness) plus every cluster size $|\mathcal{C}^{(j)}_{v}|\!\to\!\infty$ as $n\to\infty$. Consequently, under Assumption 1 and with $J\rightarrow\infty$, the test exhibits the same appealing statistical properties as the conventional KCI.
@@ 4.2 Computational Complexity
Assuming balanced partitions, each block contains roughly $n/V$ samples. The complexity of KCI per block is then $\mathcal{O}((n/V)^{3})$, and the total complexity becomes $\mathcal{O}(Jn^{3}/V^{2})$. Since each of the $J$ runs is fully parallelizable, the wall-clock cost is significantly reduced compared to the original $\mathcal{O}(n^{3})$ cost.
@@
0:  Dataset $\{(x_{i},y_{i},z_{i})\}_{i=1}^{n}$, number of partitions $V$, number of samples $J$.
@@ 5 Experiments
To empirically prove the main result in Algorithm 1, we extensively study the performance of FastKCI and compare it to the KCI implementation provided in the causal-learn package [35]. We consider different scenarios, focusing on coverage, power and causal discovery. [^4] Our full code can be found and used within the causal-learn package in python. Please see https://github.com/py-why/causal-learn/blob/main/causallearn/utils/FastKCI/FastKCI.py.
5.1 Coverage Comparison
We run experiments inspired by Zhang et al. [32], namely, we generate $n=1200$ samples of random variables $X$, $Y$ and $Z$, with $X$ and $Y$ being drawn independently conditioned on $Z$. In our scenario, we examine the type I error with growing confounding set size $D=\{1,\dots,5\}$, with all variables effecting both $X$ and $Y$ (comparable to “Case II” in Zhang et al. [32]). For the $Z_{i}$, we consider multiple ground-truth distributions, as a mixture of $V_{\text{true}}=\{1,3,10\}$ Gaussians. $X$ and $Y$ are generated as $g\left(\sum_{i}f_{i}(Z_{i})+\varepsilon\right)$ where $f$ and $g$ are random mixtures of linear, cubic and tanh functions and $\varepsilon$ is independent across $X$ and $Y$.
Figure 2: Simulated coverage of the FastKCI and the KCI at 1% and 5% levels.
@@ 5.2 Power Comparison
We repeat the experiments from above but make $X$ and $Y$ conditionally dependent by adding a small identical noise component $\nu\sim\mathcal{N}(0,\sigma^{2}_{\textrm{vio}})$ to both of them and assess the type II error. We compare KCI and FastKCI in different configurations with a growing violation of $H_{0}$ (i.e., $\sigma^{2}_{\textrm{vio}}$ is increasing). Figure 3 displays that both approaches have similar performance at a sample size of $n=1200$.
@@
For another comparison, we consider the setting from Section 5.1, but now $X$ directly causes $Y$. We calibrate the setting to a small violation, under which we observe a non-zero type II error. [^5] The magnitude of the violation is approximately $1/3$ of the signal. Table 1 shows the power under different numbers of Gaussians $V_{\text{true}}$ in the DGP.
@@ 5.3 Causal Discovery
We compare the performance of the PC algorithm using FastKCI with KCI in causal discovery tasks. For this, we consider two different settings, setting A is derived from Zhang et al. [32]. We sample $6$ random variables $\{X_{1},\dots,X_{6}\}$. For $j>i$ we sample edges with probability $0.3$. Based on the resulting DAG, we sample descendants from a Gaussian Process with mean function $\sum_{i\in\text{Pa}(X_{j})}\nu_{i}\cdot X_{i}$ (with $\nu_{i}\sim\mathcal{U}[-2,2]$) and a covariance kernel consisting of a Gaussian kernel plus a noise kernel. Setting B is derived from Liu et al. [18] and similarly consists of $6$ random variables. For $j>i$, we sample edges with probability $0.5$. The link function $f_{i}$ is randomly chosen between being linear and non-linear. For linear components, the edge weights are drawn from $\mathcal{U}[-1.5,-0.5],[0.5,1.5]$, while non-linear components follow multiple functions (sin, cos, tanh, sigmoid,polynomial). Added noise is simulated from $\mathcal{N}(0,\sigma^{2}_{i})$ with $\sigma\in\{0.2,0.5\}$. As shown in figure 4, both methods exhibit similar performance in precision, recall and F1 score of the discovered edges.
@@ 5.4 Scalability
As Figure 5 highlights, FastKCI shows excellent scalability, particularly when using a large number of blocks $V$. To further investigate the scalability of FastKCI, we scale up the sample size in the experiments and report precision, recall, f1 and computation time. [^6] To reduce complexity, in these experiments we approximate the kernel bandwidth instead of determinating it exactly with GP. See [32] for detail. The results in Table 2 show that we achieve good results in feasible time even for sample sizes where the traditional KCI is fails due to memory and CPU constraints.
@@ 5.5 Computation Time
We showcased the empirical performance of FastKCI for both pure CI tasks as well as causal discovery with PC. To shed light on the computational speed-up, we report the computation time in Figure 5), which – depending on the choice of the main tuning parameters $V$ and $J$ – is significantly faster than for the KCI.
@@ 6 Application to Production Data
We apply our proposed method to a semi-synthetic dataset for causal discovery, contained in causal-assembly [11]. The ground-truth consists of $98$ production stations, each dedicated to specific automated manufacturing processes where individual components are progressively assembled. The processes involved, such as press-in and staking, are mechanically complex and non-linear. The resulting data provides a real-world example on which causal discovery can enhance the understanding of causes for production results and yields a good benchmark for our proposed methodology.
We compare KCI with the FastKCI variant, setting $V=10$. Table 3 depicts the full assembly line with $98$ nodes, while Table 4 shows results only for one of the production stations with $16$ nodes. Please refer to Göbler et al. [11] for an overview of results with alternative CD algorithms such as PC with fisher-Z, lingam and others. In terms of precision and recall, KCI performs slightly better than FastKCI, but both outperform all tested methods in the original benchmark paper. In terms of computational speed, FastKCI, especially in the station setting with a smaller number of nodes, has a significant edge over the standard method.
@@ 7 Conclusion and Limitations
Our paper introduced FastKCI, which turns the cubic-time KCI into a more scalable and embarrassingly parallel procedure. The clustering approach in the conditioning set $Z$ via a MoE assumption is unbiased under mild conditions, while the per test cost falls roughly by a factor $1/V^{2}$. Our experiments show that FastKCI preserves KCI’s Type-I and Type-II error across diverse settings and cuts wall-clock time for both CI and causal discovery dramatically as sample size grows. Thus, our method is a promising approach for scaling up CI testing.
The main limitation is reliance on the Gaussian-mixture assumption for partitioning, although is is a common assumption for multimodal settings, misleading clusters can hurt power, particularly under small $J$. Future work should include exploring alternative sample partitioning and importance weighting schemes in order to further refine efficiency of FastKCI.
@@ Appendix A Appendix
A.1 Additional Results
@@
A.2 Kernel-based Conditional Independence Test (KCI)
Here we provide a clearly structured algorithm box summarizing the original Kernel-based Conditional Independence (KCI) test introduced by Zhang et al. [32] for clarity and comparison with the fast procedure.
Algorithm 2 Kernel-based Conditional Independence Test (KCI) by Zhang et al. [32]
@@ A.2 Kernel-based Conditional Independence Test (KCI)
6:  Bootstrap samples from the asymptotic distribution of the test statistic $\check{t}_{CI}$ [^7] For simplicity, we refer for details to Zhang et al. [32]. Sampling from the null distribution involves multiple eigendecompositions.
@@
77footnotetext: For simplicity, we refer for details to Zhang et al. [32]. Sampling from the null distribution involves multiple eigenvalue and vector computations.
Appendix B Computing Resources.
[…]

### v2

@@ 4.1 Theoretical Insight
Under the null hypothesis $H_{0}:X\perp\!\!\!\!\perp Y\mid Z$, each block statistic $T^{(j)}_{v}$ is, by exactly the same argument as in zhang2012, asymptotically a weighted sum of independent $\chi^{2}_{1}$ variables,
$T^{(j)}_{v}\;\;\xrightarrow{d}\;\;\sum_{m}\lambda_{j,v,m}\,\chi^{2}_{1,(j,v,m)},$
@@ 4.2 Computational Complexity
Assuming balanced partitions, each block contains roughly $n/V$ samples. The complexity of KCI per block is then $\mathcal{O}((n/V)^{3})$, and the total complexity becomes $\mathcal{O}(Jn^{3}/V^{2})$. Since the $J$ partitions are fully parallelizable, the wall-clock cost is significantly reduced compared to the original $\mathcal{O}(n^{3})$ cost.
While the formal complexity expression of FastKCI appears cubic in $n$, this perspective presumes that $V$ is a small, fixed constant. In practice, for the mixture model to effectively approximate the underlying distribution of the conditioning set $Z$, the number of components $V$ can increase with the sample size. This scaling ensures that the number of samples within each cluster ($n/V$) remains manageable, preventing single partitions from becoming a computational bottleneck. If $V$ is chosen to scale with $n$, the effective computational complexity of FastKCI becomes nearly linear in $n$. We support this by our experiments, where we show manageable computation times in large samples by growing $V$.
@@
0: Dataset $\{(x_{i},y_{i},z_{i})\}_{i=1}^{n}$, number of components $V$, number of partition sampling rounds $J$.
@@
4.3 Hyperparameter Selection
The practical application of FastKCI requires the specification of the hyperparameters $J$ and $V$, for which we provide some guidance. The number of partitions $J$ is primarily determined by the available computational resources. Since FastKCI is parallelized across $J$ partitions, each processed independently, increasing $J$ in line with the number of available CPU nodes does not substantially increase runtime. Moreover, because the importance sampling procedure favors partitions that more accurately model the data, a larger $J$ generally yields improved results through additional sampling rounds.
The selection of the number of clusters $V$ is comparatively more challenging. When prior knowledge exists regarding the number of Gaussian components underlying the data-generating process, such information provides a natural choice for $V$. In the more typical scenario where the distribution of the conditioning set $Z$ is unknown, selecting $V$ involves a trade-off between sample size and cluster size. Increasing the number of clusters improves the approximation of the conditional distribution of $Z$ while simultaneously reducing the average cluster size (approximately $n/V$). Smaller clusters facilitate more efficient eigendecompositions, thereby enhancing scalability. However, excessively small clusters can introduce instability. For an ablation study on $V$, see Appendix A.1444More systematic strategies for selecting $V$ have also been proposed. For example, zhang2020 recommend fitting a mixture model to the data and estimating $V$ from the posterior distribution of the mixture. Alternatively, Bayesian optimization may be employed to identify an optimal choice of $V$ in a data-driven manner.
@@ 5 Experiments
To empirically validate the main result in Algorithm 1, we extensively study the performance of FastKCI and compare it to the KCI implementation provided in the causal-learn package (zheng2024causal). We consider different scenarios, focusing on coverage, power and causal discovery [^5] A high-level implementation will be provided within a well-known causal discovery package.. Please find additional results concerning an ablation on $V$, and non-Gaussian conditioning sets in Appendix A.1.
5.1 Type-I-Error Comparison
We generate $n=1200$ samples of random variables $X$, $Y$ and $Z$, with $X$ and $Y$ being drawn independently conditioned on $Z$. In our scenario, we examine the type I error with growing confounding set size $D=\{1,\dots,5\}$, with all variables effecting both $X$ and $Y$ (comparable to “Case II” in zhang2012). For the $Z_{i}$, we consider multiple ground-truth distributions, as a mixture of $V_{\text{true}}=\{1,3,10\}$ Gaussians. $X$ and $Y$ are generated as post-nonlinear causal model $g\left(\sum_{i}f_{i}(Z_{i})+\varepsilon\right)$ where $f$ and $g$ are random mixtures of linear, cubic and tanh functions and $\varepsilon$ is independent across $X$ and $Y$.
Figure 2: Simulated Type-I-Error (“Coverage”) of the FastKCI and the KCI at 1% and 5% levels.
@@ 5.2 Power Comparison
We repeat the experiments from above, but make $X$ and $Y$ conditionally dependent by adding a small identical noise component $\nu\sim\mathcal{N}(0,\sigma^{2}_{\textrm{vio}})$ to both random variables, in order to assess the type II error. We compare KCI and FastKCI in different configurations with a growing violation of $H_{0}$ (i.e., $\sigma^{2}_{\textrm{vio}}$ is increasing). Figure 3 displays that both approaches have similar performance at a sample size of $n=1200$.
@@
For additional comparison, we consider the setting from Section 5.1, but now $X$ directly causes $Y$. We calibrate the setting to a small violation (approximately $1/3$ of the signal), under which we observe a non-zero type II error. Table 1 shows the power under different numbers of Gaussians $V_{\text{true}}$ in the DGP.
@@ 5.3 Causal Discovery
We compare the performance of the PC algorithm using FastKCI with KCI in causal discovery tasks. For this, we consider two different settings, setting A is derived from zhang2012. We sample $6$ random variables $\{X_{1},\dots,X_{6}\}$. For $j>i$ we sample edges with probability $0.3$. Based on the resulting DAG, we sample descendants from a Gaussian Process with mean function $\sum_{i\in\text{Pa}(X_{j})}\nu_{i}\cdot X_{i}$ (with $\nu_{i}\sim\mathcal{U}[-2,2]$) and a covariance kernel consisting of a Gaussian kernel plus a noise kernel. Setting B is derived from liu2024causal and similarly consists of $6$ random variables. For $j>i$, we sample edges with probability $0.5$. The link function $f_{i}$ is randomly chosen between being linear and non-linear. For linear components, the edge weights are drawn from $\mathcal{U}[-1.5,-0.5],[0.5,1.5]$, while non-linear components follow multiple functions (sin, cos, tanh, sigmoid,polynomial). Added noise is simulated from $\mathcal{N}(0,\sigma^{2}_{i})$ with $\sigma\in\{0.2,0.5\}$. As shown in Figure 4, both methods exhibit similar performance in precision, recall and F1 score between discovered edges and true causal skeleton.
@@ 5.4 Scalability
As Figure 5 highlights, FastKCI shows excellent scalability, particularly when allowing the number of components $V$ to grow with sample size, which is in accordance with our theoretical consideration. To further investigate the scalability of FastKCI, we scale up the sample size in the experiments and report precision, recall, F1 and computation time. [^6] To reduce complexity, in these experiments we approximate the kernel bandwidth instead of determinating it exactly with GP. See zhang2012 for detail. The results in Table 2 show that we achieve good results in feasible time even for sample sizes where the traditional KCI fails due to memory and CPU constraints.
@@
5.5 Comparison to Randomized Conditional Independence Test
Another recently proposed method for speeding up the KCI is the Randomized Conditional Independence Test (RCIT) (Strobl2018). The authors show that their approximation of the KCI null by random Fourier features is able to achieve significant speed-ups while being comparable to KCI in a wide range of DGPs. We demonstrate that our method that exactly replicates the KCI null law instead of approximating it, has an advantage on both type-I and type-II error when it comes to a complex, multi-modal DGP with $V=10$ Gaussians in the ground-truth. See result for type-I-error in Table 3 and for type-II in Table 8 in Appendix A.2.
FastKCI ($V=10$) KCI RCIT
$|Z|$
Type-I
Error ($\alpha=1\%$)
Type-I
Error ($\alpha=5\%$)
CPU
time [s]
Type-I
Error ($\alpha=1\%$)
Type-I
Error ($\alpha=5\%$)
CPU
time [s]
Type-I
Error ($\alpha=1\%$)
Type-I
Error ($\alpha=5\%$)
CPU
time [s]
1 0.005 0.085 3.56 0.01 0.03 19.51 0.995 1 0.016
3 0.01 0.065 6.83 0.03 0.06 118.22 0.715 0.85 0.016
5 0.005 0.04 10.55 0.03 0.06 177.66 0.265 0.39 0.016
7 0.015 0.05 22.03 0.005 0.065 163.48 0.055 0.135 0.016
10 0.005 0.045 43.02 0.01 0.02 320.21 0.03 0.11 0.016
12 0.015 0.05 63.4 0.02 0.07 323.3 0.13 0.275 0.017
15 0.01 0.04 79.31 0 0.03 669.18 0.18 0.37 0.017
30 0.005 0.035 284.63 0.005 0.05 800.66 0.35 0.54 0.018
Table 3: Type-I Errors for FastKCI, KCI and RCIT in a process with $10$ Gaussians in the ground-truth.
@@ 5.6 Computation Time
We showcased the empirical performance of FastKCI for both pure CI tasks as well as causal discovery with PC. To shed light on the computational speed-up, we report the computation time in Figure 5, which – depending on the choice of the main tuning parameters $V$ and $J$ – is significantly faster than for the KCI.
@@ 6 Application to Production Data
We apply our proposed method to a semi-synthetic dataset for causal discovery, contained in causal-assembly (goebler2024). The ground-truth consists of $98$ production stations, each dedicated to specific automated manufacturing processes where individual components are progressively assembled. The processes involved, such as press-in and staking, are mechanically complex and non-linear. The resulting data provides a real-world example on which causal discovery can enhance the understanding of causes for production results and yields a good benchmark for our proposed methodology.
We compare KCI with the FastKCI variant, setting $V=10$. Table 4 depicts the full assembly line with $98$ nodes, while Table 5 shows results only for one of the production stations with $16$ nodes. Please refer to goebler2024 for an overview of results with alternative CD algorithms such as PC with fisher-Z, lingam and others. In terms of precision and recall, KCI performs slightly better than FastKCI, but both outperform all tested methods in the original benchmark paper. In terms of computational speed, FastKCI, especially in the station setting with a smaller number of nodes, has a significant edge over the standard method.
@@ 7 Conclusion and Limitations
Our paper introduced FastKCI, which turns the cubic-time KCI into a more scalable and embarrassingly parallel procedure. The clustering approach in the conditioning set $Z$ via a mixture of experts remains statistically valid under mild conditions, while the per test cost falls roughly by a factor $1/V^{2}$. Our experiments show that FastKCI preserves KCI’s Type-I and Type-II error across diverse settings and cuts wall-clock time for both CI and causal discovery dramatically as sample size grows. Particularly, we showed empirically how FastKCI can scale causal discovery by letting the number of components $V$ grow with sample size, even under violation of the MoE assumption. If $V$ grows such that average cluster size stays constant, FastKCI approaches linearity in $n$ in computation time. Thus, our method is a promising approach for scaling up CI testing.
The main limitation is reliance on the Gaussian‑mixture assumption for partitioning, although it is a common assumption for multimodal settings and we showed FastKCIs robustness under its violation, misleading clusters can hurt power, particularly under small $J$. Further, our work does not address the issue of large conditioning set sizes $|Z|$, but it would be worth investigating a possible combination of our framwork with CI tests for large $|Z|$ (e.g., bellot2019). Future work could also include exploring alternative sample partitioning and importance weighting schemes in order to further refine efficiency of FastKCI.
8 Reproducibility Statement
We provide full source code and instructions in the supplementary material to reproduce all experiments, including simulations, and causalAssembly benchmarks. The data-generating processes for synthetic experiments are fully specified, and the semi-synthetic benchmark dataset (causal-assembly) is publicly available. Runtime environments, and computing resources are documented in Appendix B. The hyperparameters used in the experiments are specified in the according sections. The implementation of FastKCI is included and will also be released as part of an open-source causal discovery package to facilitate use by the community.
@@ Appendix A Appendix
A.1 Ablation on hyperparameter $V$
We perform an ablation study on the hyperparameter $V$. For this, we use two DGPs, one that follows Assumption 1, precisely the one from Section 5.1, with a mixture of $V=10$ Gaussians, and one, that violates it [^7] We draw $U\sim\mathrm{Unif}(-\pi,\pi)$ and set $Z=\sin(U)+0.3\sin(3U)+0.1U^{2}$. Given $Z$, we generate $X=f_{X}(Z)+\varepsilon_{X}$ and $Y=f_{Y}(Z)+\varepsilon_{Y}$, with independent noises $\varepsilon_{X},\varepsilon_{Y}\sim\mathcal{N}(0,0.5I)$, ensuring $H_{0}$ to hold while introducing post-link-noise. We use non-linear but distinct mappings (e.g., random linear projections followed by $\tanh$ or cubic terms)..
We find that the results are relatively insensitive to the choice of $V$. Small ratios of $V/n$ tend to have high computational complexity. Very large ratios, e.g. the case $V=100$ and $n=1200$ appear to not approximate the conditioning set well enough anymore and thus are not recommended. They further can increase computation time again because of inbalanced component sizes and unstable computations on small samples. Anything inbetween is recommendable.
$n=1200$ $n=5000$
$V$
Type-I
Error ($\alpha=1\%$)
Type-I
Error ($\alpha=5\%$)
CPU
time [s]
Type-I
Error ($\alpha=1\%$)
Type-I
Error ($\alpha=5\%$)
CPU
time [s]
3 0.005 0.050 48.73 0.005 0.025 1133.64
10 0.005 0.035 3.67 0.015 0.050 272.55
20 0.005 0.040 2.47 0.010 0.045 58.16
50 0.010 0.075 3.40 0.000 0.035 11.23
100 0.015 0.185 5.69 0.005 0.035 9.91
Table 6: Ablation on the choice of $V$ in FastKCI on the tests performance in a DGP with a mixture of 10 Gaussians.
n = 1200 n = 5000
$V$
Type-I
Error ($\alpha=1\%$)
Type-I
Error ($\alpha=5\%$)
CPU
time [s]
Type-I
Error ($\alpha=1\%$)
Type-I
Error ($\alpha=5\%$)
CPU
time [s]
3 0.000 0.045 45.49 0.015 0.045 1140.41
10 0.010 0.075 3.43 0.020 0.080 276.25
20 0.000 0.040 2.22 0.015 0.085 58.89
50 0.010 0.095 3.04 0.035 0.090 10.85
100 0.020 0.155 5.01 0.025 0.100 9.55
KCI 0.010 0.040 32.84
Table 7: Ablation on the choice of $V$ in FastKCI on the tests performance in a DGP with a conditioning set consisting of a non-Gaussian distribution.
Particularly, these experiments also demonstrate how FastKCIs performance appears to be not heavily reliant on Assumption 1. Both under misspecification of the number of mixture components $V$ and under a completely non-Gaussian conditioning set we observe that Type-I errors remain intact.
A.2 Power compared to RCIT
We compare the power of FastKCI to KCI and RCIT in a study similar to section 5.2. We see, that the type-II error of FastKCI is much more competitive to KCI than the one of RCIT. This further underlines our argument that in certain processes approximating the null can be insufficient.
FastKCI KCI RCIT
Violation
strength
Type-II
Error ($\alpha=5\%$)
Type-II
Error ($\alpha=1\%$)
Type-II
Error ($\alpha=5\%$)
Type-II
Error ($\alpha=1\%$)
Type-II
Error ($\alpha=5\%$)
Type-II
Error ($\alpha=1\%$)
0.10 0.960 0.840 0.905 0.795 0.975 0.915
0.15 0.835 0.680 0.505 0.340 0.940 0.810
0.20 0.325 0.175 0.155 0.080 0.690 0.480
0.25 0.015 0.010 0.015 0.000 0.380 0.190
Table 8: Type-II error comparison of KCI, FastKCI and RCIT for a process with a mixture of $10$ Gaussians. The violation strength refers to the signal of the violation relative to the signal of the influence of $Z$ on $X$ and $Y$.
A.3 Additional Results
@@
A.4 Kernel-based Conditional Independence Test (KCI)
Here we provide a clearly structured algorithm box summarizing the original Kernel-based Conditional Independence (KCI) test introduced by zhang2012 for clarity and comparison with the fast procedure.
Algorithm 2 Kernel-based Conditional Independence Test (KCI) by zhang2012
@@ A.4 Kernel-based Conditional Independence Test (KCI)
6: Bootstrap samples from the asymptotic distribution of the test statistic $\check{t}_{CI}$ [^8] For simplicity, we refer for details to zhang2012. Sampling from the null distribution involves multiple eigendecompositions.
@@
88footnotetext: For simplicity, we refer for details to zhang2012. Sampling from the null distribution involves multiple eigenvalue and vector computations.
Appendix B Implementation Details
B.1 Paritioning Scheme
In our implementation, we do not run a full EM algorithm for the Gaussian mixture. Instead, for each of the $j\in J$ partitioning rounds we use a simple empirical-Bayes sampling scheme over $Z$ approximating the NIW prior: we draw $V$ component means from a Normal centered at the empirical mean of $Z$, use an identity covariance for all components, draw mixture weights from a symmetric Dirichlet prior $\text{Dir}(\alpha)$, with $\alpha=500$ and then sample cluster assignments for each $z_{i}$ from the resulting categorical distribution (proportional to $\pi_{v}\mathcal{N}(z_{i}\mid\mu_{v},I)$). The goal of this step is to generate reasonable local partitions of the conditioning set $Z$, the importance-weighted aggregation then corrects for randomness across partitions. This lightweight scheme is sufficient in practice and avoids the extra cost of running EM inside each repetition.
B.2 Aggregation Scheme
To compute the weights $\ell^{(j,v)}$, we place independent Gaussian process regression models on the local relations $X\mid Z$ and $Y\mid Z$ within each cluster $v$. Concretely, for $i\in\mathcal{C}_{v}^{(j)}$, we assume $X_{i}=f_{X,v}(Z_{i})+\varepsilon_{X,i}$ and $Y_{i}=f_{Y,v}(Z_{i})+\varepsilon_{Y,i}$, with $f_{X,v},f_{Y,v}\sim\mathrm{GP}(0,k_{\theta})$ and Gaussian noise $\varepsilon_{\cdot,i}\sim\mathcal{N}(0,\sigma_{\cdot}^{2})$. Conditional on $Z_{\mathcal{C}_{v}^{(j)}}$ and partition indicator $U^{(j)}$, the block of observations is therefore jointly Gaussian, e.g.
$P(X_{\mathcal{C}_{v}^{(j)}}\mid Z_{\mathcal{C}_{v}^{(j)}},U^{(j)})=\mathcal{N}(0,K^{(j,v)}_{X}+\sigma_{X}^{2}I),$
and analogously for $Y$.
In the implementation, $\ell^{(j,v)}$ is computed exactly as the sum of the log-marginal likelihoods for $X\mid Z$ and $Y\mid Z$ on each block, and the weights $w_{j}$ are obtained via a softmax over the resulting partition log-likelihoods.
Appendix C Computing Resources.
The experiments in Sections 5 and 6 were performed on a high-performance computing cluster. Each node has two Intel Xeon E5-2630v3 with $8$ cores and a $2,4\textrm{GHz}$ frequency as well as $64\textrm{GB}$ RAM. For the results with a higher number of $J$, multiple nodes where used. While the runtime of single repetitions of CI or PC can be derived from Section 5.6 or respective columns in the result tables (e.g., column “Time” in Table 2), the full runtime of reproducing all experiments can be estimated around two to four weeks on a single node.
Appendix D Statement on AI Usage
For this research paper, large language models were used solely to assist with literature search, writing, and coding. All conceptualization, ideation, and theoretical contributions were carried out without AI support. The paper and code were authored entirely by the researchers, with AI serving only as a support and feedback tool.